OpenStudy (anonymous):

Solve |2y - 2| = 15 A. 8.5 B. -13 C. 8.5 and -6.5 D. 17 and -13

OpenStudy (anonymous):

it was A. 2*8.5 = 17 17-2=15

OpenStudy (anonymous):

|2y - 2| = 15 2y-2=+or-15 2y-2=15 , 2y-2=-15 2y=17 ,2y=-13 y=17/2 ,y=-13/2

OpenStudy (anonymous):

so the answer is (c)